OpenStudy (anonymous):

The goal of this problem is to solve the equation by completing the square. x^2 - 13 x + 36 = 0

OpenStudy (helder_edwin):

first add -36 to both sides

OpenStudy (anonymous):

Right, that I understand. I'm just finding it hard to think of a perfect square which would involve -13.

OpenStudy (helder_edwin):

then u got \[ \large x^2-13x+\qquad=-36 \]

OpenStudy (helder_edwin):

in the blank: put the coefficient of x (-13) halved and then squared and add the same to the right hand side

OpenStudy (anonymous):

x^2-13x+169/4=25/4

OpenStudy (anonymous):

Right?

OpenStudy (helder_edwin):

u forgot the -36 on the RHS \[ \large x^2-13x+\color{red}{\frac{169}{4}}=-36+\color{red}{\frac{169}{4}}= \frac{25}{4} \]

OpenStudy (helder_edwin):

sorry u forgot nothing.

OpenStudy (helder_edwin):

factor the LHS

OpenStudy (anonymous):

Alright. So, (x-13/2) = 25/4?

OpenStudy (anonymous):

Sorry...the LHS should be squared.

OpenStudy (helder_edwin):

yes. u r right

OpenStudy (anonymous):

Okay, wonderful. So...is the sol'n 13/2 plus or minus sqrt(25/4)?

OpenStudy (helder_edwin):

yes \[ \large x=13/2\pm\sqrt{25/4} \]

OpenStudy (anonymous):

Okay, wonderful. Thanks so much!!

OpenStudy (helder_edwin):

u r welcome